Find the vertices, eccentricity, foci and asymptotes of the hyperbola x²/8-y²/4=1
Also trace it. Under what conditions on λ the line x+λy=2 will be tangent to this hyperbola? Explain geometrically.
1
Expert's answer
2019-07-21T09:32:44-0400

We compare this equation with "x^2\/a^2-y^2\/b^2=1"


Eccentricity is "e=\sqrt{1+b^2/a^2}=\sqrt{\frac{3}{2}}"


The center is "C=(0,0)"

The vertices are "V'=(−a,0)=(−2\sqrt{2},0)" and "V=(a,0)=(2\sqrt{2},0)"

To find the foci, we need the distance from the center to the foci "c^2=a^2+b^2=12, \,\,c=\pm 2\sqrt{3}"

The foci are "F'=(-c,0)=(2\sqrt{3},0)" and "F=(c,0)=(-2\sqrt{3},0)"

The asymptotes are "x^2/8-y^2/4=0,\,\, y=\pm \frac{1}{\sqrt{2}}x"





We compare equation of tangent to hyperbola "x_0x\/a^2-y_0y\/b^2=1" with "x+\lambda y=2"

We have "x-\frac{a^2}{b^2} \frac{y_0}{x_0}y=\frac{a^2}{x_0}, \,\, x-2\frac{y_0}{x_0}y=\frac{8}{x_0}" and "x_0=4", so "y_0=\pm2" and "\lambda=-\frac{y_0}{2}=\mp 1"